A particle A has charge +q and a particle B has a charge +9q with each of them having the same mass m. If both the particles are allowed to fall from rest through the same potential difference, then the ratio of their speeds is:

Grade:12th pass

1 Answers

karthika
6 years ago
Since they are falling from rest, we have kinetic energy.Now using kinetic energy K= work done by charge=charge×potential. And also we have K=1/2mv2
Therefore, ka/kb=q×v/9q×v = 1/2mva2/1/2mvb2.
Now on solving you will get ratios of speed.